If you could, please help me with these 2 problems.
The directions says to simplify the variable expressions.
(-14)(-y^2)
and
8(-y^2)^2

Answer by mathmaven53(29)

A negative times a negative is a positive
(-14)(-y^2)=14y^2
(-y^2)^2 = (y^2)^2
= y^4
8(-y^2)^2 = 8y^4